archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    We will perform maintenance on the VDSL cards of the DSLAM LAL69 this Tuesday December 2 at midnight A loss of synchronization for few minutes is expected Date Tuesday 02 December 2014 11 06AM Reason for closing Done Comments 1

    Original URL path: http://status.ovh.es/?do=details&id=8258 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    call interruption may occur during the updating We will partially start on the night from Wednesday December 3rd to Thursday December 4th Date Monday 08 December 2014 10 46AM Reason for closing Done Comments 3 Comment by OVH Wednesday 03 December 2014 12 06PM Tonight around 4 AM we will apply the configuration change on 25 of the park Comment by OVH Thursday 04 December 2014 10 20AM The change

    Original URL path: http://status.ovh.es/?do=details&id=8246 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Comment by OVH Wednesday 26 November 2014 07 43AM The card was reset service got back running properly Comment by OVH Thursday 27 November 2014 12 50PM The default seems to still be present on the DSLAM A reboot is scheduled for 1AM tomorrow morning A down time of 15 20 minutes is expected If the problem persit an OVH technician will be sent on site tomorrow morning to diagnose

    Original URL path: http://status.ovh.es/?do=details&id=8227 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    20 November 2014 16 34PM Reason for closing Done Comments 4 Comment by OVH Thursday 20 November 2014 10 59AM We ve opened a ticket with FT Comment by OVH Thursday 20 November 2014 11 57AM An intervention by FT is underway to address a global power failure on the NRA STO93 The intervention is expected to finish at 1PM Comment by OVH Thursday 20 November 2014 16 00PM The

    Original URL path: http://status.ovh.es/?do=details&id=8201 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Details Maintenance on the fiber optics is being performed by the operator in Lille no impact for clients Date Tuesday 25 November 2014 12 59PM Reason for closing Done Comments 1 Comment by OVH Tuesday 18 November 2014 10 05AM

    Original URL path: http://status.ovh.es/?do=details&id=8179 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details Following a power outage at site SL869 the DSLAM and optical equipment at site SL869 in the city of Lyon is unresponsive Date Wednesday 12 November 2014

    Original URL path: http://status.ovh.es/?do=details&id=8133 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Task Type Modernization Category Backend Core Status CLOSED Percent Complete Details We are going to change the configuration of the DNS resolvers on the modems to distribute the load on new machines No impact is expected Date Friday 07 November 2014 19 07PM Reason for closing Done Comments 2 Comment by OVH Tuesday 04 November 2014 19 06PM The 25 of modems have the new configuration Comment by OVH Friday

    Original URL path: http://status.ovh.es/?do=details&id=8086 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Details We have a default on the equipment lil 4 n5 Right now the securing took over via lil 3 n5 Date Tuesday 04 November 2014 14 50PM Reason for closing Done Comments 1 Comment by OVH Sunday 02 November

    Original URL path: http://status.ovh.es/?do=details&id=8069 (2015-10-15)
    Open archived version from archive



  •